One number is four times another.The sum of both numbers and 12 is 37.Find the number
1
Expert's answer
2014-06-02T05:50:45-0400
Let the first number is x and the second is y. Then x=4y and x+y+12=37. Therefore we get
x=4y
x+y=25
substitute the first equation into the second
4y+y=5y=25
therefrom y=5 and x=4*5=20.
Answer: 5 and 20.